RARATCHATA found 500 ant amokars to volunteer for A study in which AACH WAR Andy A had to an ofitran traalrem UDA Group 1 recs Rescarchers found sou current smokers la valunteer for a study in which cach wiss randomly assgned to one at threr treatmenRASAAChAre found 500 tant amokers to wolunteer for a study in which AAch was randomly Aspirada na of ThiFA A YANT Groups G pResearchers found 500 current smokers la volunteer for a study in which each was randomly assigned to one of three treatmentRASRArchers tourd current smokers ba voluntar for a sh.cy in which noch was randomly nasiqnad to ana otthon treatment groups.
Answer #1

Expected Values Column 1 Column 2 Column 3 Total Row 1 64x 180 = 56.195 180 70%180 = 61.463 70x25 = 8.537 71x160 = 62.341 710

Based on the observed and expected values, the squared distances can be computed according to the following formula: (O-E)^2/E. The table with squared distances is shown below:

Squared Distances Column 1 Column 2 Column 3 Row 1 (59-61.463) 61.463 = 0.099 (62-62.341) = 0.002 62.341 (59-56.195) 56.195 -

x2 = (Oij - Eij)? Vijij Ej = 0.099 +0.711 +0.002 +0.013 +0.14 +1.008 = 1.973

Based on the information provided, the significance level is \alpha = 0.05α=0.05 , the number of degrees of freedom isdf=(2−1)×(3−1)=2, so then the rejection region for this test is R={χ2:χ2>5.991}.

táto = 1.97

Xéri = 5.991

The test statistic is less than critical value therefore there is NOT sufficient evidence .

f) GROTEDT 54 DID NOT QUIT QUIT GROUP 1 GROUP 2 GROUP 3 OPTION B

STAY CONSTANT

MOST SMOKERS DO NOT WANT TO QUIT

G] THERE IS NOT SUFFICIENT EVIDENCE TO SHOW THAT MOST SUBJECT WANT TO QUIT.